[ Japanese | English ]

COLD NEWS

The Nobeyama 45-m Radio Telescope Web site was renewed on 1st August 2018. Some pages deleted are removed during this renewal. If there is important information for you without a link, please consult via the Helpdesk.

Pre-announcement of Call for Proposals for the next observing season (2017-2018) is made. (2017-06-15)
The deadlines, observing categories, etc. are available. The details will be infomred in future.
Announcemnent of Call for Proposals for Development for Nobeyama 45-m Telescope. (2017-05-12)
An update on the Suspension of the Open Use Observation. (2017-04-17)
Announcemnent of Suspension of the Open Use Observation. (2017-03-30)
Recent TA* of SiO masers is updated. (2017-03-14)
Examples for OTF observations with the FOREST receiver are illusterated. (2017-02-06)
Rules for backup observations are updated. (2017-02-06)
Imformation for a limitation in the Device Table is updated. (2017-02-01)
Please follow the instruction when you observe in the 2016-2017 observing season.
Instruction to make observing scripts for the FOREST receiver is updated. (2017-01-19)
If you observe with the FOREST receiver in this observing season, please follow this procedure.
Procedures for receiver tuning is updated. (2017-01-15)
Recent skyline around the Nobeyama 45-m telescope is updated. Trees at Az=0-90deg. are pruned.(2016-12-13)
Proposal Information for the open-use observation from December 2016 to May 2017 is open. (2016-06-21)
Antenna efffciencies for lower frequency during 2015-2016 observing season are added. (2016-06-12)
A bug of frequency shift in the position switching data was found during 2012-2016 observation period. Details are shown in the link. (2016-05-02)
Status of Continuum Backend is updated. #11 is unavailable. (2016-03-21)
A List of Sample files for Receiver Setting page is linked from Device Tab page about Obstable Z (2016-01-19)
Observation manual is updated (2015-12-03)
A bug fix version of Time Estimator for OTF is released. (2015-11-30)
Please re-estimate observing time if you who are going to propose OTF observations.
Proposal Deadline of General 2nd Period is extended to 7th December from 1st December due to a bug of Time Estimator. (2015-11-27)
The release date of Time Estimator (bug fix version) will be 30th November
Pages for Obstable Z are updated (2015-11-17)
A guide after an acceptance of your proposal until observations is added (2015-11-17)
A figure of requested hours in the 1st period of the 2015-2016 season at each LST is appended in Call for Proposals page (2015-11-13)
Frequently Asked Questions on submitting a proposal are summrizrd on A Guide for Proposal Submission page (2015-11-02)
Update the skyline of the 45-m telescope in "How to Plan an Observation using the 45-m Telescope" page (2015-09-04)
Proposal Information - Call for Proposals (2015-08-03)
Antenna efffciency 2014-2015 (2015-05-26)
Helpdesk is open (2015-03-06)
Information on the new 4-beam receiver FOREST (2014-08-14)
Proposal Information - Call for Proposals (2014-07-31)
Spectra of Standard Sources, 2013 - (2014-04-25)
New spectrometer SAM45 status
Antenna efffciency 2013-2014 (2014-04-17)
S80 receiver has a trouble at present. (2014-02-05)
AOS-W3 has a trouble at present. (2014-02-04)
One of the AD converters (PANDA) in the new system has been replaced to an improved one, and it is connected to IF 13 and 14. (Generally PANDA produces spurious lines, but the improved PANDA should not produce such lines.) (2013-12-24)
Accepted proposal (2013-10-16)
Proposal Information - Call for Proposals (2013-07-24)
Notification of closing server operation for data analysis for NRO users (2013-05-16)
Antenna efffciency 2012-2013 (2013-04-12)
At present, one of the beams of BEARS ch3-5 cannot be used with the 1st local frequency of higher than 108 GHz. Please take this situation into account for your planning. ( 12CO cannot be observed with this beam, but 13CO can be observed with the USB setting.) (2013-02-27)
The T100 receiver (2SB) at the 3 mm region was decommissioned. (2013-01-08)
Instead of T100, please use the TZ1 receiver with similar specifications. TZ1 can be connected to the new spectrometer SAM45. We are revising the related web pages. Please use the software Obstable Z (zobs) for preparing the observing script of TZ1.
A method to prepare device tables in Obstable Z (zobs) has been changed.(2013-01-08)
Information on the new 4-beam receiver FOREST (2012-08-29)
AOS status (2012-03-02)
CASA reduction manual ( in Japanese )(2011/12/16)
Spectra of Standard Sources is here (2011-01-14)
Java version of Newstar (reduction software for spectral lines) was released.
In addition, a task (beta version) of Newstar "New Spectrum FITS" (to transform spectral files to FITS format) was also released.
For details click here. (2010-01-06)
How to start OBSTABLE for 45m is here.
The proposal cover sheet has been upgraded. Please use ver 4.10.(2008-10-30)
Proposal information for open-use observation in 2009 season is now open.(2008-09-03)
The deadline of General Program has been postponed to 11 November 2008.
The LO of H30 receiver does not work stably in some frequency. Please ask us for details. (2008-05-20)
IF1 attenuator is broken (May 6-13, 2008)
The 1dB-step attenuator of IF #1 is broken. Please do not use IF 1.
Export version of NOSTAR
Export version of OTF reduction software is released. (Mar. 31, 2008)
Antenna efficiencies for 2007-2008 season
The antenna efficiencies in the 2007-2008 season is presented. (Mar. 13, 2008)
A bandpass filter of IF2 cannot be used. (Feb. 25-26, 2008)
The 5.1--5.6 GHz bandpass filter of the IF 2 was broken, and seems to be "spontaneously" repaired.
AOS-H7 cannot be used. (Dec. 29, 2007-Feb. 5, 2008)
The LO attenuator of the AOS-H6 is broken.
AOS-H6 cannot be used. (Jan. 3-8, 2008)
The step attenuator of the AOS-H6 sometimes does not function.
Export (beta) version of NOSTAR
Public beta release of Export version of OTF reduction software. (2007-11-30)
BEARS star formation project
Data from BEARS star formation project is open. (2007-08-28)
Status report
Status Report for 2007-2008 season is now open.(2007-06-15)
Proposal information
Proposal information for open-use observation in 2007-2008 season is now open. (2007-06-06)
IF7 attenuator is broken (Apr. 27-May 1, 2007)
The 1dB-step attenuator of IF #7 is broken. Please do not use IF 7.
Beam efficiencies for 2006-2007 season
Beam efficiencies are released. (2007-04-25)
BEARS scaling factors for 2006-2007 season
BEARS calibration data (scaling factors) are released. (2007-04-19)
BEARS (CH15) cannot be used. (Mar. 19-21, 2007)
Due to a snapping of an IF cable, CH15 of BEARS cannot be used since Mar. 19. And a trouble of the refrigerator which occurred succeedingly makes all the 25 beams unavailable (Mar. 20). Both troubles are fixed by Mar. 21.
The AOS-H7 cannot be used (Mar. 1-13, 2007)
The 4th local oscillator for AOS-H7 is broken.
IF2 attenuator is broken (Feb. 22-27, 2007)
The 1dB-step attenuator of IF #2 is broken. Please do not use IF 2.
Trouble on AC-03 (Feb. 13-14, 2007)
Due to a trouble on a digital spectrometer A03, CH13 of BEARS cannot be used.
We have renewed the Web site. (Feb. 7, 2007)
Old page (Japanese)
Status of AOS-W1 (Feb. 6-14, 2007)
The band characteristics of AOS-W1 was found to be bad. Please do not assign important lines only in this spectrometer.
--> It turned out to be OK.
Operation resumed
The 45 m telescope was temporary repaired, and the operation (common use observations) has started again. For details see the top page of this observatory. (Feb. 5, 2007)
Operation Interrupted (Jan.-Feb. 5, 2007)
The operation of the 45 m telescope is now stopped due to the trouble of the driving system. For details see the top page of this observatory. (Jan. 24, 2007)
Schedule of open use observations in 2006-2007 season
Open use will start from early Jan. 2007. Now we are preparing for it. (Dec. 24, 2006)
[ collimator: done | subref: done | pointing: done | beam squint: done | beam efficiencies: almost done | etc. ]
We now open the OTF observing mode.
OTF (On The Fly) mode is available for open use.
Change of 45-m manager
N. Kuno has became the manager of the telescope in succession to K. Sunada. (Nov. 2006)
The frequency converter of AOS-H3 is now broken.
Unfortunitely AOS-H3 cannot be used. (March 28 - April 4, 2006)
The H40 receiver was replaced.
The 1st IF and the side band were changed (9.8 ---> 6.0 GHz, LSB ---> USB). You can get a sample file from the window of the device table.
Present status of AOS (Dec. 2004-)
Present status of BEARS
At present a channel 25 cannot be used due to the breaking of the wire (Dec. 31, 2004-Jan. 27, 2005).
The intensity calibration system of the old port (old MMC) is broken (Jan. 8-19, 2005)
Please use LOAD1 (LD1) instead of the old MMC for the observations with the H40, H32/28, and H22 receivers.
For this coming observational season, we are measuring the telescope performance and adjusting the basic parameters (subreflector, pointing, beam squint, efficiency, etc.).
The adjustment of the subreflector position and the revision of pointing parameters have already finished.
The operation of the 45m radiotelescope in this season finished. (31 May 2004-)
On 27 May the operation in this observational season finished. In addition to the usual maintenance works, we are going to replace the control system of the telescope in this summer.
Trouble of AOS-W4 (13 April 2004-)
The baseline of AOS-W4 may be bad. Please do not assign important lines only in this spectrometer.
The revisions of OBSTABLE (2 Dec. 2003-)
The use of the digital backend for the single beam receivers (2 Dec. 2003-)
From this observing season, it is possible to connect the digital backend to the single beam receivers. The manual will be available soon. For details please ask to Dr. Sunada.
Present status of AOS (Dec. 2003-)
Trouble of AOS-W4 (9-13 April 2004)
At present fringe noise appears at half of the channel ranges of AOS-W4. Please do not assign important lines only in this spectrometer.
Trouble of AOS-H8 (2 Dec. 2003 - 9 March 2004)
At present the final local oscillator for AOS-H8 is broken, and consequently, H8 cannot be used.
Trouble of an attenuator at IF7 6-2 GHz converter (8-9 Dec. 2003)
Please do not use the IF7, or use it with even number (dB) of attenuation in the device table (ask to the assistant). In this case, please do not assign important lines only in the IF7.
Now the 45m telescope is not operated for the maintenance. (deleted 28 Nov. 2003)
Problem of the band selection board (1 April 2003 -)
The S100 receiver cannot be connected to the IF 7 line. Please do not assign the IF 7 line for the S100 receiver.
PLL problem of the S100 receiver (18 March 2003 -)
The PLL of S100 does not work well at the local frequency of 87.925 GHz. Please avoid to use this local frequency.
Tuning of the S100/S80 receivers (Dec. 2002 -)
Almost complete tuning can be done by PC at the control room. In this tuning process bias voltage is not adjusted. If this adjustment is necessary to obtain maximum performance, please ask to the assistant.
Present status of BEARS
Information on the local oscillators of the H22 receiver
Please note that the tuning ranges of the local (Gunn) oscillators of H22 is not completely continuous. The tuning ranges are 14.75-15.25 GHz, 16.7-17.25 GHz, and 18.7-19.3 GHz. Their will be no big problem, because the bandwidth of H22 is rather wide (2 GHz).
Present status of AOS (Dec. 2002-
Connection of cables for continuum observations (NOBA and other receivers)(Dec. 2002-)
This connection became automatic by using a switch box. Please click 'start tuning' in the main console to make this connection effective before continuum observations.
Low performance of the S100 receiver (ca. 11 March 2003 -) (deleted 1 April 2003)
The performance of S100 is unfortunitely slightly low at present. Roughly speaking it is better to use S80 instead of S100 for observations below 100 GHz. For observations above 100 GHz S100 seems better than S80. For details please ask the people maintaining the receivers.
Trouble of IF6 (19 March 2003 -) (deleted 25 March 2003)
The system temperature of AOSs connected to IF6 is now abnormally low. Please avoid to use IF6.
Trouble of an attenuator (14 March 2003-) (deleted 17 March 2003)
An attenuator of IF6 at 6 to 2 GHz converter is now broken. Please do not use the IF6 line.
Present status of BEARS (deleted 18 Feb. 2003)
Trouble of an attenuator at IF8 6-2 GHz converter (7-17 Feb. 2003)
In case of odd number (dB) of attenuation, signal does not pass. Please do not use the IF8, or use it with even number (dB) of attenuation (ask to the assistant).

2018-08-01 Update